# Fermi-liquid transport beyond the upper critical field in superconducting La2PrNi2O7 thin films

### v1

@@
Yu-Te Hsu***These authors contributed equally to this work.ythsu@phys.nthu.edu.tw Department of Physics, National Tsing Hua University, Hsinchu 300044, Taiwan Department of Materials Science and Engineering, National Tsing Hua University, Hsinchu 300044, Taiwan Yidi Liu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Physics, Stanford University, Stanford, CA 94305, United States Yoshimitsu Kohama Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Tommy Kotte Dresden High Magnetic Field Laboratory (HLD-EMFL), Helmholtz-Zentrum Dresden-Rossendorf, Dresden 01328, Germany Vikash Sharma Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Yaoju Tarn Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Yijun Yuyuyijun@stanford.edu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Harold Y. Hwang Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States
@@
Unconventional superconductivity typically emerges out of a strongly correlated normal state, manifesting as a Fermi liquid with highly enhanced effective mass or a strange metal with $T$-linear resistivity in the zero-temperature limit. In Ruddlesden-Popper bilayer nickelates $R_{3}$i2O7, superconductivity with a critical temperature ($T_{\rm c}$) exceeding 80 and 40 K has been respectively realised in bulk crystals under high pressure and thin films under compressive strain. These advancements create new materials platforms to study the nature of high-$T_{\rm c}$ superconductivity, calling for the characterisation of fundamental normal-state and superconducting parameters therein. Here we report detailed magnetotransport experiments on superconducting La2PrNi2O7 (LPNO) thin films under pulsed magnetic fields up to 64 T and access the normal-state behaviour over a wide temperature range between 1.5 and 300 K. We find that the normal state of LPNO exhibits the hallmarks of Fermi liquid transport, including $T^{2}$ temperature dependence of resistivity and Hall angle, and $H^{2}$ magnetoresistance obeying Kohler scaling. Using the empirical Kadowaki-Woods ratio relating the transport coefficient and electronic specific heat, we estimate a quasiparticle effective mass $m^{*}/m_{e}\simeq 10$ in PLNO, thereby revealing the highly renormalized Fermi liquid state which hosts the high-temperature nickelate superconductivity. Our results demonstrate that LPNO follows the same $T_{\rm c}/T_{\rm F}^{*}$ scaling observed across a wide variety of strongly correlated superconductors and provide crucial constraints for a viable model for superconductivity in bilayer nickelates.
@@ Introduction
The recent discovery of superconductivity at 80 K in (La,Pr)3Ni2O7 under high pressure makes the bilayer Ruddlesden-Popper (RP) nickelates the latest material system exhibiting high-temperature superconductivity (HTSC) [1, 2]. Despite intense interests in the nature of superconductivity in RP bilayer nickelates [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], key parameters characterising the normal and superconducting state are still missing. This is in large part due to the high-pressure condition required to induce superconductivity, which limits the type of experiments that can be performed to primarily transport under static magnetic field and x-ray diffraction. The subsequent realisation of superconductivity in epitaxially strained (La,Pr)3Ni2O7 thin films with a critical temperature $T_{\rm c}\approx$ 40 K (refs. [13, 14, 15]) creates a new material platform to which a broader selection of experimental probes can be employed at ambient pressure, including photoemission [16, 17], x-ray spectroscopy [18], and electron energy-loss spectroscopy [19]. Moreover, the availability of superconducting (La,Pr)3Ni2O7 thin films enables a straightforward implementation of pulsed high-field magnetotransport experiment, which can reveal a wealth of information on its superconducting and normal state, including upper critical field ($H_{\rm c2}$), carrier concentration ($n_{\rm H}$), and functional form of low-temperature normal-state resistivity $\rho_{\rm n}(T)$.
@@
Prior magnetotransport experiments on LPNO thin film estimate that $\mu_{0}H_{\rm c2}$ in the $T=0$ limit is well above 50 T [13, 14, 15], necessitating the use of pulsed field conditions to access the low-$T$ normal state. Here, we report detailed magnetotransport experiment on superconducting LPNO thin films under pulsed magnetic fields up to 64 T. We reveal salient characteristics of the normal-state transport in LPNO, including 1) a $T^{2}$ resistivity in the field-induced low-$T$ normal state, 2) a saturating Hall coefficient as $T\rightarrow 0$ and a $T^{2}$ temperature dependence of the Hall angle, and 3) an observation of Kohler scaling in normal-state magnetoresistance. These findings provide strong evidence that the underlying electronic ground state in strained LPNO thin film is a Fermi liquid. Our measurements also reveal that $H_{\rm c2}(T=0)$ in LPNO exhibits an anisotropy with $\gamma=H_{c2,\parallel}/H_{\rm c2,\perp}\approx 2.5$. By estimating the magnitude of key normal-state and superconducting parameters using established approaches, we find that the ratio between $T_{\rm c}$ and effective Fermi temperature $T_{\rm F}^{*}$ in bilayer RP nickelates exhibits a scaling of $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$, as found in a wide variety of strongly-correlated superconductors [21, 22, 23, 24], thereby hinting at a universal underlying principle governing the $T_{\rm c}$ magnitude in unconventional superconductors.
@@ Results
Figure 1a shows the in-plane resistivity $\rho_{\rm xx}(H,T)$ of a LPNO thin film in pulsed magnetic field up to 53 T. The field is applied parallel to the crystalline $c$-axis i.e. perpendicular to the film surface. A clear transition from a superconducting to a resistive state can be seen below $T_{\rm c}=41$ K (defined using the 90% $\rho_{\rm n}$ criterion with $\rho_{\rm n}(T)$ obtained by the parallel-resistor fit to resistivity above 50 K). The temperature dependence of resistivity measured in the field-induced normal state at 53 T and extrapolated to 0 T are plotted in Fig. 1b. A prominent feature of $\rho_{\rm xx}(T)$ is the positive curvature as $T\rightarrow 0$ and negative curvature as $T\rightarrow 300$ K. As demonstrated in previous work [15], $\rho_{\rm xx}(T)$ above $T_{\rm c}$ can be well-described using a parallel resistor formula (PRF):
@@
where $\rho_{0}$, $\rho_{\rm max}$, and $A_{2}$ are fitting constants. By suppressing superconductivity with large magnetic field, we find that the field-induced normal-state resistivity below $T_{\rm c}$ tracks the PRF fit very closely and shows a clear saturation towards an apparent $T^{2}$-behaviour as $T\rightarrow 0$ (Fig. 1b inset). The same behaviour is reproduced in a second sample (Supplementary Materials Fig. S1). Furthermore, the extrapolated zero-field resistivity exhibits essentially the same $\rho(T)$ form as $\rho_{\rm xx}(\textrm{53\leavevmode\nobreak\ T})$ (see details for zero-field resistivity extrapolation in Methods), demonstrating that the normal-state $\rho_{\rm xx}(T)$ of LPNO exhibits a $T^{2}$ functional form in the $T=0$ limit, characteristic of a Fermi-liquid ground state.
Figure 1: In-plane resistivity of a superconducting La2PrNi2O7 thin film. a, Magnetoresistivity isotherms $\rho_{\rm xx}(H)$ up to 53.4 T measured at the following temperatures: 100, 80, 60, 50, 45, 40, 35, 30, 25, 20, 15, 10, 7.0, 4.1, and 1.5 K. b, Temperature-dependent resistivity $\rho_{\rm xx}(T)$ at 0 and 53 T. Grey dash line is a fit to the measured zero-field resistivity above between 50 and 300 K using a parallel resistor model: $1/\rho(T)=1/(\rho_{0}+A_{2}T^{2})+1/\rho_{\rm max}$, which finds $\rho_{0}=102\leavevmode\nobreak\ {\rm\mu\Omega\leavevmode\nobreak\ cm}$, $\rho_{\rm max}=778\leavevmode\nobreak\ {\rm\mu\Omega\leavevmode\nobreak\ cm}$, and $A_{2}=8.1\leavevmode\nobreak\ \rm{n\Omega\leavevmode\nobreak\ \leavevmode\nobreak\ cm\leavevmode\nobreak\ K^{-2}}$, respectively. Inset: $\rho_{\rm xx}$ versus $T^{2}$ below 60 K, showing a $\Delta\rho(T)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ T^{2}$ behaviour in the measured 53-T resistivity and extrapolated zero-field resistivity.
@@
Figure 2a shows the Hall resistivity $\rho_{\rm yx}(H)$ measured between 10 and 60 K on the same sample as shown in Fig. 1. $H$-linear behaviour in $\rho_{\rm yx}(H)$ (within the noise level) is found in the (field-induced) normal state, consistent with the low-field measurements above $T_{\rm c}$ in La3Ni2O7 and LPNO thin films [13, 15]. The temperature dependence of Hall coefficient $R_{\rm H}(T)$ between 10 and 240 K, plotted in Fig. 2a inset, shows a monotonic increase as $T$ decreases and approaches a constant value of $R_{\rm H}\approx$ 0.3 mm3/C as $T\rightarrow 0$. We estimate the total carrier density using $n_{\rm H}(0)\simeq 1/(eR_{\rm H}(0))=2.1\times 10^{22}\leavevmode\nobreak\ \rm{cm}^{-3}$ and from which a two-dimensional (2D) carrier density $n_{\rm 2D}(0)=n_{\rm H}(0)d=2.1\times 10^{15}\rm\leavevmode\nobreak\ {cm}^{-2}$ per NiO2 bilayer unit ($d=10.3$ is the distance between neighboring NiO2 bilayers [2]). Note that this $n_{\rm H}$ value should be interpreted as a coarse estimate for the inference of $k_{\rm F}$ and possibly not constitute an accurate quantification of the total carrier density, as details of the electronic structure in bilayer nickelates remain to be established.[25, 16] Nevertheless, assuming a free-electron model in two dimensions, we find $k_{\rm F}=1.15\times 10^{10}\leavevmode\nobreak\ \rm{m}^{-1}$, comparable to that of the $\beta$ sheet found by recent photoemission experiments[25, 16].
@@
Figure 2: Hall resistivity, Kohler scaling, and Hall angle in La2PrNi2O7 thin film. a, Hall resistivity isotherms $\rho_{\rm yx}(H)$ measured at indicated temperatures. Dash lines are fits made to the normal-state $\rho_{\rm yx}$ at high fields using $R_{\rm H}=\rho_{\rm yx}/(\mu_{0}H)$. Inset: Hall coefficient versus temperature $R_{\rm H}(T)$. A saturation of $R_{\rm H}\simeq 0.3$ mm3/C is found as $T\rightarrow 0$. Vertical dashed line marks $T_{\rm c}$. b, Normalised magnetoresistance versus magnetic field scaled by zero-field resistivity i.e. $\rho_{\rm xx}/\rho_{\rm xx}(0)$ versus $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler plot. Normal-state data measured at indicated temperatures collapses into a single curve following $\rho_{\rm xx}/\rho_{\rm xx}(0)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ (\mu_{0}H/\rho_{\rm xx}(0))^{2}$ shown by the dashed line. c, Inverse Hall angle $\cot\theta_{\rm H}=\rho_{\rm xx}/\rho_{\rm xx}$ versus $T^{2}$. Filled points correspond to measurements at 53 T and open points correspond to extrapolated 53 T values using Hall resistivity measured at 8 T (i.e. $\rho_{\rm yx}(\rm 53\leavevmode\nobreak\ T)$ = $R_{\rm H}(8\leavevmode\nobreak\ \rm{T})\times(53/8)$). Dashed line is a fit using $\cot\theta_{\rm H}=C+BT^{2}.$
@@
The form of normal-state magnetoresistance (MR = $\Delta\rho_{\rm xx}(H)/\rho_{\rm xx}(0)$) is examined in Fig. 2b. The MR magnitude is small ($\lesssim 4\%$ up to 53 T) and follows a $H^{2}$-behaviour i.e. $\Delta\rho(H)/\rho(0)\leavevmode\nobreak\ \propto\leavevmode\nobreak\ H^{2}$. We find that $\rho_{\rm xx}(H)/\rho_{\rm xx}(0)$ collapse into a single curve when plotted against $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler scaling, which implies that the MR magnitude is dictated by the carrier mean free path $\ell\sim\omega_{\rm c}\tau$ (i.e. the product of cyclotron frequency $\omega_{\rm c}$ and carrier lifetime $\tau$), as expected for a Fermi-liquid state characterised by a single quasiparticle lifetime [26]. The fact that the normal-state MR $\propto\leavevmode\nobreak\ H^{2}$ and obeys Kohler scaling allows us to extract the zero-field normal-state resistivity below $T_{\rm c}$ as shown in Fig. 1b (see Methods for details). The Fermi-liquid nature of magnetotransport in LPNO is further supported by the observation of a negligible MR when the magnetic field is applied parallel to the sample surface i.e. $\mathbf{H}\parallel ab$ (Supplementary Materials Fig. S2) and a $T^{2}$ dependence of the Hall angle: $\cot\theta_{\rm H}=\rho_{\rm xx}/\rho_{\rm yx}\leavevmode\nobreak\ \propto\leavevmode\nobreak\ (\omega_{\rm c}\tau)^{-1}$ (Fig. 2c). These features are in stark contrast to the high-$T_{\rm c}$ cuprates within the strange-metal regime, in which a sizeable MR is found with $\mathbf{H}\parallel ab$ (ref. [27]) and the longitudinal resistivity and Hall angle exhibit distinct $T$-dependence [28], suggesting that the longitudinal and Hall transport are governed by two distinct carrier lifetimes. Overall, we find that the normal-state transport characteristics of superconducting LPNO, including $\rho_{\rm xx}(T)$, MR, Hall effect, and Hall angle, can all be well-described using the standard transport theory for a conventional Fermi-liquid, the key finding of this work.
@@
Next we extract the upper critical field $H_{\rm c2}$ in both $\mathbf{H}\parallel c$ and $\mathbf{H}\parallel ab$ configuration (i.e. $H_{\rm c2,\perp}$ and $H_{\rm c2,\parallel}$, respectively). Due to the considerable transition width in magnetic field, the choice of different extraction criteria leads to different forms of $H_{\rm c2}(T)$ (see Supplementary Materials Fig. S3). We adopt the 90% $\rho_{\rm n}$ criterion, in accordance with previous reports [13, 15], and extract $H_{\rm c2}$ using the condition $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}(\rm{53\leavevmode\nobreak\ T)}=0.9$. The $\mu_{0}H_{\rm c2}(T)$ data is fitted using the linearised Ginzburg-Landau form [29]:
@@
where $\phi_{0}$ is the flux quantum, $\xi_{ab}(0)$ is the in-plane coherence length at zero temperature, and $d_{\rm sc}$ is the effective superconducting thickness. A highly anisotropic $H_{\rm c2}$ is found with $\mu_{0}H_{\rm c2,\perp}(0)=(42.8\pm 0.4)$ T and $\mu_{0}H_{\rm c2,\parallel}(0)=(106\pm 1)$ T, and a corresponding anisotropic factor $\gamma=H_{\rm c2,\parallel}/H_{\rm c2,\perp}=2.48$. Furthermore, we find the in-plane coherence length $\xi_{ab}(0)=(2.76\pm 0.01)$ nm and a superconducting thickness $d_{\rm sc}=(3.88\pm 0.02)$ nm, comparable to the film thickness of 5 nm and suggesting a bulk nature of superconductivity.

### v2

@@
Yu-Te Hsu***These authors contributed equally to this work.ythsu@phys.nthu.edu.tw Department of Physics, National Tsing Hua University, Hsinchu 300044, Taiwan Center for Quantum Science and Technology, National Tsing Hua University, Hsinchu 300044, Taiwan Department of Materials Science and Engineering, National Tsing Hua University, Hsinchu 300044, Taiwan Yidi Liu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Physics, Stanford University, Stanford, CA 94305, United States Yoshimitsu Kohama Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Tommy Kotte Dresden High Magnetic Field Laboratory (HLD-EMFL), Helmholtz-Zentrum Dresden-Rossendorf, Dresden 01328, Germany Vikash Sharma Institute for Solid State Physics, The University of Tokyo, Kashiwa, Chiba 277-8581, Japan Yaoju Tarn Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Bai Yang Wang Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Zhi-Xun Shen Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Geballe Laboratory for Advanced Materials, Department of Physics and Applied Physics, Stanford University, Stanford, CA 94305, United States Stanford Synchrotron Radiation Lightsource, SLAC National Accelerator Laboratory, Menlo Park, CA 94025, United States Yijun Yuyuyijun@stanford.edu Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States Harold Y. Hwang Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, Menlo Park, Stanford, CA 94025, United States Department of Applied Physics, Stanford University, Stanford, CA 94305, United States
@@
Unconventional superconductivity typically emerges out of a strongly correlated normal state, manifesting as a Fermi liquid with highly enhanced effective mass or a strange metal with $T$-linear resistivity in the zero-temperature limit. In Ruddlesden-Popper bilayer nickelates $R_{3}$Ni2O7, superconductivity with a critical temperature ($T_{\rm c}$) exceeding 80 and 40 K has been respectively realised in bulk crystals under high pressure and thin films under compressive strain. These advancements create new materials platforms to study the nature of high-$T_{\rm c}$ superconductivity, calling for the characterisation of fundamental normal-state and superconducting parameters therein. Here we report detailed magnetotransport experiments on superconducting La2PrNi2O7 (LPNO) thin films under pulsed magnetic fields up to 64 T and access the normal-state behaviour over a wide temperature range between 1.5 and 300 K. We find that the normal state of thin-film LPNO exhibits the hallmarks of Fermi-liquid transport, including $T^{2}$ temperature dependence of resistivity and Hall angle, and $H^{2}$ magnetoresistance obeying Kohler scaling. Using the empirical Kadowaki-Woods ratio relating the transport coefficient and electronic specific heat, we estimate a quasiparticle effective mass $m^{*}/m_{e}\simeq 10$ in thin-film LPNO, thereby revealing the highly renormalized Fermi liquid state which hosts the high-temperature nickelate superconductivity. Our results demonstrate that thin-film LPNO follows the same $T_{\rm c}/T_{\rm F}$ scaling observed across a wide variety of strongly correlated superconductors and establish key characteristics of the normal ground state from which the superconductivity in epitaxially strained bilayer nickelates emerges.
@@ Introduction
The recent discovery of superconductivity at 80 K in (La,Pr)3Ni2O7 under high pressure makes the bilayer Ruddlesden-Popper (RP) nickelates the latest material system exhibiting high-temperature superconductivity [2, 3]. Despite intense interests in the nature of superconductivity in RP bilayer nickelates [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], key parameters characterising the normal and superconducting state are still missing. This is in large part due to the high-pressure condition required to induce superconductivity, which limits the type of experiments that can be performed to primarily transport under static magnetic field and x-ray diffraction. The subsequent realisation of superconductivity in epitaxially strained (La,Pr)3Ni2O7 thin films with a critical temperature $T_{\rm c}\approx$ 40 K (refs. [14, 15, 16]) creates a new material platform to which a broader selection of experimental probes can be employed at ambient pressure, including photoemission [17, 18], x-ray spectroscopy [19], and electron energy-loss spectroscopy [20]. Moreover, the availability of superconducting (La,Pr)3Ni2O7 thin films enables a straightforward implementation of pulsed high-field magnetotransport experiment, which can reveal key information on its superconducting and normal state, including upper critical field ($H_{\rm c2}$), scaling laws in magnetotransport, and functional form of low-temperature normal-state resistivity $\rho_{\rm n}(T)$.
@@
Prior magnetotransport experiments on LPNO thin films estimate that $\mu_{0}H_{\rm c2}$ in the $T=0$ limit is well above 50 T [14, 15, 16], necessitating the use of pulsed field conditions to access the low-$T$ normal state. Here, we report detailed magnetotransport experiment on superconducting LPNO thin films under pulsed magnetic fields up to 64 T. We reveal salient characteristics of the normal-state transport in LPNO, including 1) a $T^{2}$ resistivity in the field-induced low-$T$ normal state, 2) a $T^{2}$ temperature dependence of the Hall angle, and 3) an observation of Kohler scaling in normal-state magnetoresistance. These findings provide strong evidence that the underlying electronic ground state in compressively strained LPNO thin film is a Fermi liquid. Our measurements also reveal that $H_{\rm c2}(T=0)$ in LPNO thin film exhibits an anisotropy with $\gamma\approx 2.5$ , measured by the ratio between $H_{\rm c2}$ with field applied parallel ($H_{\rm c2,\parallel}$) and perpendicular to the $a$-$b$ plane ($H_{\rm c2,\perp}$), respectively. By estimating the magnitude of key normal-state and superconducting parameters using established approaches, we find that the ratio between $T_{\rm c}$ and effective Fermi temperature $T_{\rm F}$ in bilayer RP nickelates exhibits a scaling of $T_{\rm c}/T_{\rm F}\sim 0.01$, as found in a wide variety of strongly-correlated superconductors [22, 23, 24, 25], thereby hinting at a universal underlying principle governing the $T_{\rm c}$ magnitude in unconventional superconductors.
@@ Results
Figure 1a shows the in-plane resistivity $\rho_{\rm xx}(H,T)$ of a LPNO thin film under pulsed magnetic field up to 53 T. The field is applied parallel to the crystalline $c$-axis i.e. perpendicular to the film surface. A clear transition from a superconducting to a resistive state can be seen below $T_{\rm c}=41$ K (defined as the temperature at which the measured $\rho_{\rm xx}(T)$ falls below $0.9\rho_{\rm n}(T)$ with $\rho_{\rm n}(T)$ obtained by the parallel-resistor fit to the normal-state resistivity; Fig. 1b). The temperature dependence of resistivity measured in the field-induced normal state at 53 T and extrapolated to 0 T are plotted in Fig. 1b. A prominent feature of $\rho_{\rm xx}(T)$ is the positive curvature as $T\rightarrow 0$ and negative curvature as $T\rightarrow 300$ K. As demonstrated in previous work [16], $\rho_{\rm xx}(T)$ above $T_{\rm c}$ can be well-described using a parallel resistor formula (PRF)[26]:
@@
where $\rho_{0}$, $\rho_{\rm max}$, and $A_{2}$ are fitting constants (For discussion of the physical significance of $\rho_{\rm max}$, refer to Supplementary Materials Sec. E). By suppressing superconductivity with large magnetic field, we find that the field-induced normal-state resistivity below $T_{\rm c}$ closely tracks the PRF fit and shows a clear saturation towards an apparent $T^{2}$-behaviour as $T\rightarrow 0$ (Fig. 1b inset). The same behaviour is reproduced in a second sample (Supplementary Materials Sec. I). Furthermore, the extrapolated zero-field resistivity exhibits essentially the same $\rho(T)$ form as $\rho_{\rm xx}(\textrm{53\penalty 10000\ T})$ (see details for zero-field resistivity extrapolation in Methods and Supplementary Materials Sec. D), demonstrating that the normal-state $\rho_{\rm xx}(T)$ of LPNO exhibits a $T^{2}$ functional form in the $T=0$ limit, characteristic of a Fermi-liquid ground state.
Figure 1: In-plane resistivity of a superconducting La2PrNi2O7 thin film. a, Magnetoresistivity isotherms $\rho_{\rm xx}(H)$ up to 53.4 T measured at the following temperatures: 100, 80, 60, 50, 45, 40, 35, 30, 25, 20, 15, 10, 7.0, 4.1, and 1.5 K. b, Temperature-dependent resistivity $\rho_{\rm xx}(T)$ measured at 0 T (solid line) and 53 T (filled circles), and the extrapolated 0-T values (open circles). Grey dash line is a fit to the measured zero-field resistivity above between 50 and 300 K using a parallel resistor model: $1/\rho(T)=1/(\rho_{0}+A_{2}T^{2})+1/\rho_{\rm max}$, which finds $\rho_{0}=102\penalty 10000\ {\rm\mu\Omega\penalty 10000\ cm}$, $\rho_{\rm max}=780\penalty 10000\ {\rm\mu\Omega\penalty 10000\ cm}$, and $A_{2}=8.1\penalty 10000\ \rm{n\Omega\penalty 10000\ cm\penalty 10000\ K^{-2}}$, respectively. Inset: $\rho_{\rm xx}$ versus $T^{2}$ below 60 K, showing a $\Delta\rho(T)\penalty 10000\ \propto\penalty 10000\ T^{2}$ behaviour in the measured 53-T resistivity and extrapolated zero-field resistivity. For a discussion of the possible origin and impact of two-step transition below $T_{\rm c}$, refer to Supplementary Materials Sec. C.
@@
Figure 2a shows the Hall resistivity isotherms $\rho_{\rm yx}(H,T)$ measured between 10 and 60 K on the same sample as shown in Fig. 1. $H$-linear behaviour in $\rho_{\rm yx}(H)$ (within the noise level) is found in the (field-induced) normal state, consistent with the low-field measurements above $T_{\rm c}$ in La3Ni2O7 and LPNO thin films [14, 16]. The magnitude of Hall coefficient $R_{\rm H}(T)$ between 10 and 240 K, plotted in Fig. 2a inset, shows a monotonic increase as $T$ decreases and $R_{\rm H}$ approaches 0.3-0.3 mm3/C as $T\rightarrow 10$ K. Since the Fermi surface of thin-film LPNO consists of a small electron-like $\alpha$-pocket and a large hole-like $\beta$-pocket [27, 17, 28], the negative sign of $R_{\rm H}$ likely reflects the multiband nature of its electrical transport[29], thereby precludes the inference of normal-state carrier density using the measured $R_{\rm H}$, as similarly found in the infinite-layer nickelates [30].
@@
Figure 2: Hall resistivity, Kohler scaling, and Hall angle in La2PrNi2O7 thin film. a, Hall resistivity isotherms $\rho_{\rm yx}(H)$ measured at indicated temperatures. Traces at different temperatures are shifted successively by 1 $\mu\Omega$ cm for clarity. Dash lines are fits made to the normal-state $\rho_{\rm yx}$ at high fields using $R_{\rm H}=\rho_{\rm yx}/(\mu_{0}H)$. Inset: Hall coefficient versus temperature $R_{\rm H}(T)$. $R_{\rm H}$ approaches 0.3-0.3 mm3/C as $T\rightarrow 0$. Vertical dashed line marks $T_{\rm c}$. b, Normalised magnetoresistance versus magnetic field scaled by zero-field resistivity i.e. $\rho_{\rm xx}/\rho_{\rm xx}(0)$ versus $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler plot. Normal-state data measured at indicated temperatures collapses into a single curve following $\rho_{\rm xx}/\rho_{\rm xx}(0)\penalty 10000\ \propto\penalty 10000\ (\mu_{0}H/\rho_{\rm xx}(0))^{2}$ shown by the dashed line. Note that traces of $T$ = 10, 20, 30, and 40 K deviate from the Kohler scaling function at low $\mu_{0}H/\rho_{\rm xx}(0)$ until normal-state behaviour is recovered at sufficiently high fields. $\rho_{\rm xx}$ ($\rho_{\rm yx}$) traces measured below $T_{\rm c}$ are (anti-)symmetrized, whereas the 50- and 60-K data are measured using the positive-polarity trace only, which nonetheless show good agreement with measurement performed at 14 T. c, Inverse Hall angle $\cot\theta_{\rm H}=|\rho_{\rm xx}/\rho_{\rm yx}|$ versus $T^{2}$. Filled points correspond to measurements at 53 T and open points correspond to extrapolated 53 T values using Hall resistivity measured at 8 T (i.e. $\rho_{\rm yx}(\rm 53\penalty 10000\ T)$ = $R_{\rm H}(8\penalty 10000\ \rm{T})\times(53/8)$). Dashed line is a fit using $\cot\theta_{\rm H}=C+BT^{2}.$
@@
The form of normal-state magnetoresistance (MR = $\rho_{\rm xx}(H)-\rho_{\rm xx}(0)]/\rho_{\rm xx}(0)$) is examined in Fig. 2b. The MR magnitude is small ($\lesssim 4\%$ up to 53 T) and follows a $H^{2}$-behaviour i.e. $\Delta\rho(H)/\rho(0)\penalty 10000\ \propto\penalty 10000\ H^{2}$. We find that $\rho_{\rm xx}(H)/\rho_{\rm xx}(0)$ collapse into a single curve when plotted against $\mu_{0}H/\rho_{\rm xx}(0)$, known as the Kohler scaling, implying that the MR magnitude is dictated by the carrier mean free path $\ell\sim\omega_{\rm c}\tau$ (i.e. the product of cyclotron frequency $\omega_{\rm c}$ and carrier lifetime $\tau$), as expected for a Fermi-liquid state characterised by a single quasiparticle lifetime [31]. The fact that the normal-state MR $\propto\penalty 10000\ H^{2}$ and obeys Kohler scaling allows us to extract the zero-field normal-state resistivity below $T_{\rm c}$ as shown in Fig. 1b (see Methods and Supplementary Sec. D for extraction details). The Fermi-liquid nature of magnetotransport in thin-film LPNO is further supported by the observation of a negligible MR when the magnetic field is applied parallel to the sample surface i.e. $\mathbf{H}\parallel ab$ (Supplementary Materials Fig. S5) and a $T^{2}$ dependence of the Hall angle: $\cot\theta_{\rm H}=|\rho_{\rm xx}/\rho_{\rm yx}|\penalty 10000\ \propto\penalty 10000\ (\omega_{\rm c}\tau)^{-1}$ (Fig. 2c). These features are in stark contrast to the high-$T_{\rm c}$ cuprates within the strange-metal regime, in which a sizeable MR is found with $\mathbf{H}\parallel ab$ (ref. [32]) and the longitudinal resistivity and Hall angle exhibit distinct $T$-dependence [33], suggesting that the longitudinal and Hall transport are governed by two distinct carrier lifetimes. Overall, we find that the normal-state transport characteristics of superconducting LPNO thin film, including $\rho_{\rm xx}(T)$, MR, and Hall angle, can all be well-described using the standard transport theory for a conventional Fermi-liquid, the key finding of this work.
@@
Next we extract the upper critical field $H_{\rm c2}$ in both $\mathbf{H}\parallel c$ and $\mathbf{H}\parallel ab$ configuration (i.e. $H_{\rm c2,\perp}$ and $H_{\rm c2,\parallel}$, respectively). Due to the considerable transition width in magnetic field, the choice of different extraction criteria leads to different forms of $H_{\rm c2}(T)$ (Supplementary Materials Fig. S7). We adopt the 90% $\rho_{\rm n}$ criterion, in accordance with previous reports [14, 16], and extract $H_{\rm c2}$ using the condition $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}(\rm{53\penalty 10000\ T)}=0.9$. The $\mu_{0}H_{\rm c2}(T)$ data, shown in Fig. 3, is fitted using the linearised Ginzburg-Landau form [34]:
@@
where $\phi_{0}$ is the flux quantum, $\xi_{ab}(0)$ is the in-plane coherence length at zero temperature, and $d_{\rm sc}$ is the effective superconducting thickness. An anisotropic $H_{\rm c2}$ is found with $\mu_{0}H_{\rm c2,\perp}(0)=(42.8\pm 0.4)$ T and $\mu_{0}H_{\rm c2,\parallel}(0)=(106\pm 1)$ T, corresponding to an anisotropic factor $\gamma=H_{\rm c2,\parallel}/H_{\rm c2,\perp}=2.48$. We further find that the $\mu_{0}H_{\rm c2}(\theta\simeq 90^{\circ})$ data measured at 35 K can be well fitted using the two-dimensional (2D) Tinkham model [35] (Fig. 3 inset), pointing to a 2D nature of superconductivity (see Supplementary Sec. F for details of $H_{\rm c2}$ extraction and model descriptions). This $\gamma$ value is comparable to optimally doped YBa2Cu3O7-δ ($\gamma\approx 2.2$; ref. [36]) and larger than infinite-layer nickelate La0.8Sr0.2NiO2 ($\gamma\approx 1.7$; ref. [37]) and iron-based superconductor Ba(Fe,Co)2As2 ($\gamma\approx 1.1$; ref. [38]), implying that the anisotropy in electronic structure of thin-film LPNO is similar to the high-$T_{\rm c}$ cuprates. Furthermore, we find the in-plane coherence length $\xi_{ab}(0)=(2.76\pm 0.01)$ nm and a superconducting thickness $d_{\rm sc}=(3.88\pm 0.02)$ nm, comparable to the film thickness of 5 nm and suggesting a bulk nature of superconductivity.